# Gujarat-382355, India

### v1

@@ 5.2 Both Field-Field and Momentum-Field Coupling Enabled
When we turn on both the field-field and the momentum-field couplings, we observe a superposition of the characteristics of the individual couplings discussed in the previous subsection. For example, linear entropy shows an early recoherent phase, followed by a late-time saturation indicating a transition to a maximally-mixed state (see figs. 4(a) and 5(a)). One can see that the COF faithfully reproduces this behaviour. It shows a bump at early times – followed by a plateau – which signifies the recoherence. At late-times, the time evolution of COF shows a sharp kink, followed by a linear growth, which signals decoherence (see figs. 4(b) and 5(b)).
The physical reason behind the dynamical behaviour of the linear entropy, and the COF, can be understood as follows. As long as the field-field coupling $\lambda$ is not zero, it always dominates at late-times and leads to complete decoherence of the system quantum field. This can be understood by examining the strength of the field-field coupling as compared to that of the field-momentum coupling $\rho$. The former is accompanied by a factor of $a^{2}$ while the latter by that of $a$, and since the scale factor grows exponentially during inflation, the $\lambda^{2}a^{2}$ term wins over the $\rho a$ term (see eq. 8) if one waits long enough. The second feature to note is that the linear entropy, corresponding to smaller momentum-field coupling, crosses over that for a larger value of $\rho$ during this late-time phase once decoherence takes over [^4] From fig. 5(a), this can be seen as the orange curve crosses over the green and red curves at late-times.. A similar behavior can be seen for the COFs of the corresponding plots in fig. 5. This is because the faster decoherence (or growth in $S_{l}$ or COF) will take place for the smaller value of $\rho$ as this term tries to slow down this behavior. In addition, we notice for a fixed $\rho$, the stronger field-field coupling $\lambda$ shifts the timescale of decoherence to an earlier time, whereas the recoherence timescale remains constant (see fig. 4).
It is important to note and compare the timescales of the features that correspond to the recoherence and to the saturation of decoherence for the linear entropy and the COF time evolution as demonstrated in fig. 6. A comparison between the timescales of these characteristic features from the dynamical behavior of the linear entropy and that of the COF shows a clear similarity (see fig. 7). This, again, indicates the recoherence and the decoherence signatures of the evolution of the COF suggested in section 5.1. In fig. 7, the COF recoherence timescale is identified at the point where the COF temporarily “flattens out” and remains roughly constant before the later phase of decoherence. Similarly, the recoherence timescale in linear entropy is marked by the point at which the linear entropy returns to 0 after the initial temporary phase of decoherence. Finally, the decoherence timescales are characterized by the late-time “bump” in COF (as discussed in Bhattacharyya_2024 ) and the point at which the linear entropy is sufficiently close to saturation at 1.555Since the linear entropy never reaches exactly 1, we chose 0.998 as a “sufficiently close” benchmark.
@@ Acknowledgment
AB is supported by the Core Research Grant (CRG/2023/ 001120) and Mathematical Research Impact Centric Support Grant (MTR/2021/000490) by the Department of Science and Technology Science and Engineering Research Board, India. AB also acknowledges the associateship program of the Indian Academy of Sciences (IASc), Bengaluru. SB is supported in part by the Higgs Fellowship and by the STFC Consolidated Grant “Particle Physics at the Higgs Centre”. SSH is supported in part by the “Quantum Technologies for Sustainable Development” grant from the National Institute for Theoretical and Computational Sciences of South Africa (NITHECS).

### v2

@@ 5.2 Both Field-Field and Momentum-Field Coupling Enabled
When we turn on both the field-field and the momentum-field couplings, we observe a superposition of the characteristics of the individual couplings discussed in the previous subsection. For example, linear entropy shows an early recoherent phase, followed by a late-time saturation indicating a transition to a maximally-mixed state (see Fig. (4(a)) and Fig. (5(a))). One can see that the COF faithfully reproduces this behaviour. It shows a bump at early times – followed by a plateau – which signifies the recoherence. At late-times, the time evolution of COF shows a sharp kink, followed by a linear growth, which signals decoherence (see Fig. (4(b)) and Fig. (5(b))).
@@
At this point, it is important to compare the timescales of these above characteristic features corresponding to the recoherence and the saturation of decoherence for the linear entropy and the COF time evolution—previously illustrated separately in the left and right panels of Figs. (4) and (5). Figure (6) overlays the dynamics of both quantities for two representative parameter choices, clearly demonstrating the alignment of their characteristic timescales. In addition, over a range of coupling strength parameters $\lambda$ and $\rho$, the timescales of these characteristic features from the linear entropy and the COF evolution are compared in Fig. (7) further revealing a clear similarity. This reinforces the interpretation of these features as signatures of recoherence and decoherence in the COF dynamics suggested in Sec. (5.1). More precisely, the COF recoherence timescale—plotted in Fig. (7)—is identified at the point where the COF temporarily “flattens out” and remains roughly constant before the later phase of decoherence. Similarly, the recoherence timescale in linear entropy is marked by the point at which the linear entropy returns to 0 after the initial temporary phase of decoherence. Finally, the decoherence timescales are characterized by the late-time “bump” in COF (as discussed in Bhattacharyya_2024 ) and the point at which the linear entropy is sufficiently close to saturation at 1.777Since the linear entropy never reaches exactly 1, we chose 0.998 as a “sufficiently close” benchmark.
The physical reason behind the dynamical behaviour of the linear entropy, and the COF, can be understood as follows. As long as the field-field coupling $\lambda$ is not zero, it always dominates at late-times and leads to complete decoherence of the system quantum field. This can be understood by examining the strength of the field-field coupling as compared to that of the field-momentum coupling $\rho$. Firstly, the former is accompanied by a factor of $a^{2}$ while the latter by that of $a$, and recall that the scale factor grows exponentially during inflation. More importantly, the form of the operators corresponding to the two terms is such that the $\zeta^{\prime}\mathcal{F}$ term is proportional to the decaying mode of the adiabatic perturbation, which decays very fast after horizon crossing. This is the main reason why the field-field coupling term wins over the field-momentum coupling term (see eq. 8) if one waits long enough. The second feature to note is that the linear entropy, corresponding to smaller momentum-field coupling, crosses over that for a larger value of $\rho$ during this late-time phase once decoherence takes over [^8] From Fig. (5(a)), this can be seen as the orange curve crosses over the green and red curves at late-times.. A similar behavior can be seen for the COFs of the corresponding plots in Fig. (5). This is because the faster decoherence (or growth in $S_{l}$ or COF) will take place for the smaller value of $\rho$ as this term tries to slow down this behavior. In addition, we notice for a fixed $\rho$, the stronger field-field coupling $\lambda$ shifts the timescale of decoherence to an earlier time, whereas the recoherence timescale remains constant (see Fig. (4)).
@@ Acknowledgment
AB is supported by the Core Research Grant (CRG/2023/ 001120) and Mathematical Research Impact Centric Support Grant (MTR/2021/000490) by the Department of Science and Technology Science and Anusandhan National Research Foundation (formerly SERB), Government of India. AB also acknowledges the associateship program of the Indian Academy of Sciences (IASc), Bengaluru. SB is supported in part by the Higgs Fellowship and by the STFC Consolidated Grant “Particle Physics at the Higgs Centre”. SSH is supported in part by the “Quantum Technologies for Sustainable Development” grant from the National Institute for Theoretical and Computational Sciences of South Africa (NITHECS).